There are 12 coins that consist of quarters and dimes. The value of these coins is $2.25
Write a system of equations:
Solve the system of equations:
How many quarters and dimes do I have?

Found 2 solutions by Alan3354, jorel1380:
Answer by Alan3354(69443)

d + q = 12
10d + 25q = 225
--------------
You can do the rest, right?

Answer by jorel1380(3719)

x+y=12
.1x+.25y=2.25
-----------------
Multiply the second equation by 100:
10x+25y=225
10x+10y=120
15y=105
y=7
x=5
There are 5 dimes and 7 quarters.